# Single-Shot Readout and Weak Measurement of a Tin-Vacancy Qubit in Diamond

### v1

@@ Appendix B Extended discussion
B.1 Understanding the ${}^{-}$ Hamiltonian
The ${}^{-}$ Hamiltonian is predicted by ab initio calculations in Ref. [40]. In our previous work, Ref. [26], these calculations are compared to experimental values obtained by fitting the eigenvalues of the ${}^{-}$ Hamiltonian to measurement of its level structure vs. magnetic field orientation. There are many parameters in this fit including strain, spin-orbit coupling, the orbital Zeeman effect, anisotropy of the Zeeman effect, and the vector orientation of the spin dipole. This fit is under-determined if every parameter of the ${}^{-}$ Hamiltonian is taken as a free parameter, and so to constrain the fit we made assumptions: (1) The direction normal to the surface of the chip is parallel to the $\hat{z}$-axis of the magnetic coil. (2) The magnet calibration is accurate and magnet hysteresis is negligible. (3) Higher order terms in the ${}^{-}$ Hamiltonian are neglected [40].
The values of the ${}^{-}$ Hamiltonian reported in Table 1 of Ref. [26] rely on these assumptions, which are imperfect at some level. While this analysis was a more precise determination of the ${}^{-}$ Hamiltonian than prior experimental work [20, 21, 23, 22], it can be improved upon by future experiments that do not rely on these assumptions. For example, subsequent work in Ref. [24] reports slightly different values of the ${}^{-}$ Hamiltonian but constrains the fit by studying multiple emitters, by using a three-axis vector magnet to optimize alignment with the spin dipole, and by quantifying the magnet’s calibration.
@@ B.2 Estimating angular misalignment
Understanding the ${}^{-}$ Hamiltonian helps to understand spin polarization and readout, including by determining alignment between the magnetic field and spin dipole moment. In our current experiment we use a two-axis vector magnet, only, which prohibits arbitrary alignment between the magnet and spin. To estimate the minimum misalignment we compare the measurement of cylicity vs. $\zeta$ (Fig. 2e) to a model based on the driven ${}^{-}$ Hamiltonian (see Appendix B.3 of Ref. [26]), and using the Hamiltonian parameters from Ref. [24]. This model closely follows our data for an azimuthal misalignment of $\varphi\approx 10\degree$. Use of the alternate Hamiltonian parameters reported in Ref. [26] has minimal effect on this result.
B.3 Comparison to prior work
@@
Reference
Qubit
fidelity, $F_{r}$
counts, $\bar{n}_{b}$
duration, $\tau$
Device
Spin Control?
Robledo et al. [42]
93%
6.4
$40\,\mathrm{\mu s}$
SIL
Yes
Sukachev et al. [12]
89%
6.2
$20\,\mathrm{ms}$
Bulk
Yes
Gorlitz et al. [39]
74%
1.2
$200\,\mathrm{\mu s}$
Bulk
No
Parker et al. [32]
${}^{-}$ (nuclear)
80%
1.6
$15\,\mathrm{\mu s}$
Waveguide
No
This work
87.4%
4.3
$50\,\mathrm{\mu s}$
Mesa
Yes
Despite subtleties in determining the ${}^{-}$ Hamiltonian and angular misalignment, it is clear that the readout fidelity reported in this work uses an ${}^{-}$ center which is only somewhat aligned with the external $\vec{B}$ field, has a spin polarization time of $\approx 20\,\mathrm{\mu s}$, and has a cyclicity of $\approx 2\times 10^{3}$. At this operating condition we achieve single shot readout of the ${}^{-}$ center’s electronic spin with an average of up to $\approx 4$ photons collected during a $50\,\mathrm{\mu s}$ integration window. In Table 1, we compare these results to prior examples of single-shot spin readout of color centers in diamond.
Looking to the future, the misalignment and modest collection efficiency of our experiment emphasize how much ${}^{-}$ readout can be improved. Even with these non-idealities, the rate of photon collection in our work is much greater than that in the first demonstration of single-shot readout of an ${}^{-}$ electron spin [12], which reported an average of $\approx 6$ photons collected during a 20 ms readout window. This may reflect the ${}^{-}$’s higher quantum efficiency, $\eta_{q}$: Ref. [40] calculates $\eta_{q}=0.14$ for the ${}^{-}$ and $\eta_{q}=0.91$ for the ${}^{-}$. Subsequent work on ${}^{-}$ readout has shown greater collection efficiency over faster timescales by optimized use of nanophotonics [15, 16, 17]. Our results here suggest such improvement will also be possible using the ${}^{-}$.
We also compare our results to Ref. [39], to our knowledge the only prior published single-shot readout of an ${}^{-}$ electronic spin. Ref. [39] demonstrates readout of $74\%$ fidelity (max of $1.2$ mean counts per shot) collected during a $200\,\mathrm{\mu s}$ readout window. This result is achieved in a bulk sample on an unstrained ${}^{-}$. Lack of strain yields high-cyclicity with reduced sensitivity to alignment of the magnetic field, but also suppresses the magnetic dipole transition necessary for spin control [26]. Finally, we note that Ref. [32] measures single-shot readout of a nuclear spin via a waveguide-integrated ${}^{-}$, obtaining a single-shot nuclear spin readout fidelity of $80\%$ in $15\,\mathrm{\mu s}$. In comparison, our work demonstrates a regime in which both single-shot readout of an ${}^{-}$’s electronic spin, and, rapid microwave control of this spin are possible.
@@ C.1 Setup
The setup used in this experiment, Fig. 8, is similar to that in Ref. [26] but optimized for higher transmission on the collection path. The repump (532 nm) and resonant (M-Squared) lasers are both modulated by acousto-optic modulator (4C2C-532-AOM and 4C2C-633-AOM, Gooch and Housego). Typically, we operate with the excitation laser detuned by several GHz from the ${}^{-}$’s optical transitions, and drive these transitions using the first sideband created by an electro-optic modulator (PM-0S5-PFU-PFU-620, Eospace). By combining two microwave signal generators (SG396, SRS) using a microwave power combiner, we can drive the A1 transition only, the B2 transition only, or both simultaneously as is necessary for implementing charge resonance checks. Pulsing is controlled by an arbitrary waveform generator (Pulse Streamer 8/2, Swabian).
For spin control, microwaves are delivered by a wire-bond draped over the diamond chip as in Ref. [26], and using a similar microwave power ($48\,\mathrm{dBm}$ into the cryostat). The wire-bond has been replaced from that used in Ref. [26]; we attribute the slower Rabi rate here to the new wire-bond simply being further from the spin.
To focus excitation light and collect signal confocally, a cryogenic objective (LT-APO/VISIR/0.82, Attocube) is placed directly above the chip. This objective has a focal length of 2.87 mm, a numerical aperture of 0.82, and a working distance of 0.65 mm. Signal is then routed out of the cryostat to a 90:10 beamsplitter, whose 90% port leads to a free-space to fiber coupler. This coupler is connected by a multimode optical fiber (FG025LJA, core diameter $25\pm 3\,\mathrm{\mu m}$, Thorlabs) to a single photon counting module (SPCM-AQR-15-FC, Perkins Elmer). Detection events are time-tagged (Picoharp300, PicoQuant).
Before the fiber coupler we place a 635 nm long pass filter (635 nm Longwave EdgeBasic, Semrock) to collect the ${}^{-}$’s phonon sideband (PSB), only, thus filtering most excitation scatter and also the ${}^{-}$’s zero-phonon line. Note that this wavelength still filters a fraction of the PSB; a different filter closer to the zero phonon line would have lead to greater counts, e.g. a filter at 625 nm. Finally, a 532 nm notch filter (532 nm StopLine, Semrock) is also placed on the collection path to mitigate noise from the repump laser; even though this laser is toggled off during readout, there is some non-zero feed through of the AOM used to toggle the green laser which can scatter into the collection path, and so this notch filter is still useful.
@@ C.2 Sources of loss
Source
Efficiency
Reference / Method
Path to Improve
Quantum efficiency: $\eta_{q}$
$80\%-90\%$
[20, 40]
@@
Debye-Waller factor: $\eta_{\mathrm{psb}}$
$43\%$
[23]
resonant collection (cross-polarization)
@@
Scattering from chip: $\eta_{\mathrm{phot}}$
$5\%$
simulation, Fig. 9
nanopillars, waveguides, cavity-QED
Optical path loss: $\eta_{\mathrm{path}}$
$35\%$
transmission, at 670 nm
optimized optics
Detector: $\eta_{\mathrm{det}}$
$65\%$
datasheet spec., at 650 nm
superconducting detectors
Total: $\eta$
$0.4\%$
$\eta=\eta_{q}\eta_{\mathrm{psb}}\eta_{\mathrm{phot}}\eta_{\mathrm{path}}\eta_{\mathrm{det}}$
all of the above
@@
Figure 9: (a) Device. The $\langle 100\rangle$ diamond chip contains arrays of nanophotonic structures, both pillars and mesas. Structures have a height of $1\,\mathrm{\mu m}$ and Sn atoms are implanted at a depth of $\approx 90\pm 20\,\mathrm{nm}$ [26]. (b) Scan to optimize focal position. The center of the measured lobe (which originates from imaging with a multi-mode fiber under slightly de-focused conditions) corresponds to focal alignment at the edge of a mesa. A scan range of 0.2 V corresponds to a lateral distance of $\approx 1\,\mathrm{\mu m}$. (c) Numerical simulation of the far-field scattering pattern from an ${}^{-}$ center in a mesa. (d) Numerical simulation of 645 nm light collected from both a mesa and (e) a pillar. Simulations include the fraction of light scattered upward, the modal overlap of the objective and the scattering pattern, and their product which gives the overall fraction of light routed to the collection path. From the simulation in (d) we conclude $\eta_{\mathrm{phot}}\approx 5\%$ of emission enters the collection path. In comparison, similar simulations show a scattering efficiency of $3\%$ from an emitter in bulk diamond at a depth of 100 nm.
@@
In Table 2, we estimate the sources of inefficiency and optical loss expected for this work. Multiplying all sources of inefficiency yields: $\eta=\eta_{q}\eta_{\mathrm{psb}}\eta_{\mathrm{phot}}\eta_{\mathrm{col}}\eta_{\mathrm{det}}\approx 0.4\%$. This qualitatively matches the measured values of $\eta\approx 0.2\%$ (Fig. 3) and $\eta\approx 0.1\%$ (Fig. 4).
@@
The dominant source of loss is internal reflection of light into the diamond sample, Fig. 9. By fine tuning the emitter position using piezoelectric nanopositioners, we find this emitter to actually be in a mesa structure, rather than a nanopillar as stated in Ref. [26]. For an ${}^{-}$ center implanted at a depth of 100 nm and for structures $1\,\mathrm{\mu m}$ in height, simulations determine scattering from bulk diamond to be $\approx 3\%$, scattering from a mesa is $\approx 5\%-9\%$, Fig. 9d, and scattering from a pillar is $\approx 7\%-28\%$, Fig. 9e. These values are reported at 645 nm, near the maximum of the ${}^{-}$’s phonon sideband (PSB). Given the proximity of the color center to a pillar, we use values for an ${}^{-}$ at the edge of the mesa structure and estimate the (average across the entire PSB) fraction $\eta_{\mathrm{phot}}\approx 5\%$ of emitted light is scattered into our collection path. This number accounts for the modal mismatch between the far field for the ${}^{-}$ emission and the Gaussian mode accepted by the cryo-objective.
@@
Other sources of loss include the emitter’s quantum efficiency (the probability of non-radiative decay), filtering of the ${}^{-}$’s ZPL, loss in the optical path, and detector inefficiency. Finally, Table 2 may not capture every source of loss. Note that we use a 635 nm long pass filter to collect the PSB (see Appendix C.1 for details); this removes $\approx 14.5\%$ of the PSB light, and a cutoff of $\approx 624\,\mathrm{nm}$ would be best to avoid extra loss [20, 21, 22, 23]. Additionally, while we use multi-mode fiber to maximize photon collection, some imperfection in the coupling could still be present.
@@
Optimized diamond nanophotonics provides a clear path toward high improvement in collection efficiency (e.g. $\approx 25\%$ for inverse-designed grating couplers [49], and $\approx 90\%$ for tapered fibers combined with 1D photonic crystal cavities [50]). These devices have also been demonstrated in conjunction with stable ${}^{-}$’s [28, 29, 30, 32, 34, 33]. The other sources of loss can be improved upon too. For example, loss associated with filtering of the resonant excitation laser can be avoided by filtering by polarization rather than by wavelength (routinely done in the form of cross-polarization detection). Optical loss in the collection path can be mitigated with higher quality optics, e.g. by using higher reflectivity mirrors, and by replacing the 90:10 beamsplitter on the collection path with a more transmissive component. Detector inefficiency can be improved by switching to state-of-the-art superconducting detectors.
@@
With improved nanophotonics only, we expect a collection efficiency of order ×$\times 10$ or more, and with all possible improvements we expect an efficiency of order ×$\times 100$ or more can be achieved in future experiments. Even modest improvements in efficiency will allow for much higher single-shot readout fidelity (since fidelity is non-linear with number of photons), operation at lower drive power (see Fig. 4c), faster readout, and, single-shot readout at external magnetic fields with greater misalignment.

### v2

@@ Appendix B Extended discussion
B.1 Understanding the SnV- Hamiltonian
The SnV- Hamiltonian is predicted by ab initio calculations in Ref. [44]. In our previous work, Ref. [30], these calculations are compared to experimental values obtained by fitting the eigenvalues of the SnV- Hamiltonian to measurement of its level structure vs magnetic field orientation. There are many parameters in this fit including strain, spin-orbit coupling, the orbital Zeeman effect, anisotropy of the Zeeman effect, and the vector orientation of the spin dipole. This fit is underdetermined if every parameter of the SnV- Hamiltonian is taken as a free parameter, and so to constrain the fit we made assumptions: (1) The direction normal to the surface of the chip is parallel to the $\hat{z}$ axis of the magnetic coil. (2) The magnet calibration is accurate and magnet hysteresis is negligible. (3) Higher-order terms in the SnV- Hamiltonian are neglected [44].
The values of the SnV- Hamiltonian reported in Table 1 in Ref. [30] rely on these assumptions, which are imperfect at some level. While this analysis was a more precise determination of the SnV- Hamiltonian than prior experimental work [11, 25, 27, 26], it can be improved upon by future experiments that do not rely on these assumptions. For example, subsequent work in Ref. [28] reports slightly different values of the SnV- Hamiltonian but constrains the fit by studying multiple emitters, by using a three-axis vector magnet to optimize alignment with the spin dipole, and by quantifying the magnet’s calibration.
@@ B.2 Estimating angular misalignment
Understanding the SnV- Hamiltonian helps to understand spin polarization and readout, including by determining alignment between the magnetic field and spin dipole moment. In our current experiment we use a two-axis vector magnet, only, which prohibits arbitrary alignment between the magnet and spin. To estimate the minimum misalignment we compare the measurement of cylicity vs $\zeta$ (Fig. 2e) to a model based on the driven SnV- Hamiltonian (see Appendix B.3 in Ref. [30]) and using the Hamiltonian parameters from Ref. [28]. This model closely follows our data for an azimuthal misalignment of $\varphi\approx 10\degree$. Use of the alternate Hamiltonian parameters reported in Ref. [30] has minimal effect on this result.
B.3 Importance of strain and magnetic field alignment
[…]
Figure 9: Strain considerations for cyclicity and spin control. X axis: the ratio of the variable ground-state strain $2\Upsilon_{g}$, to the fixed ground-state spin-orbit coupling $\lambda_{g}/2\pi=822\,\mathrm{GHz}$ [28]. Strain $2\Upsilon_{g}$ is defined such that the ground-state splitting is $\sqrt{\lambda_{g}^{2}+(2\Upsilon_{g})^{2}}$. Left y axis (blue): cyclicity (solid lines) for dc magnetic field misalignment of $0.1\degree$, $1\degree$, and $10\degree$. Right y -axis (red): Rabi frequency when driven by an ac microwave field perpendicular to the spin dipole moment (dashed lines). Results fall on top of each other for field misalignment of $0.1\degree$, $1\degree$, and $10\degree$.
The simulations in Fig. 9 illustrate a strain-dependent relationship between spin control and readout. Decreased strain increases cyclicity, which improves readout to the detriment of spin control. Increased strain improves spin control but to the detriment of lower cyclicity for a misaligned magnetic field. Cyclicity increases as the magnetic field becomes more aligned; this slows the spin polarization time via resonant driving (e.g., Fig. 2d) but is helpful for spin readout. Since field alignment has negligible effect of microwave Rabi frequency, it is desirable to work in a highly aligned field, in parallel to previous SiV- experiments [15].
[…]
B.4 Comparison to prior work
@@
Reference Qubit Fidelity, $F_{r}$ Counts, $\bar{n}_{b}$ Duration, $\tau$ Device Spin control?
Robledo et al. [46] NV- 93% 6.4 $40\,\mathrm{\mu s}$ SIL Yes
Sukachev et al. [15] SiV- 89% 6.2 $20\,\mathrm{ms}$ Bulk Yes
Görlitz et al. [43] SnV- 74% 1.2 $200\,\mathrm{\mu s}$ Bulk No
Parker et al. [36] SnV- (nuclear) 80% 1.6 $15\,\mathrm{\mu s}$ Waveguide No
This work SnV- 87.4% 4.3 $50\,\mathrm{\mu s}$ Mesa Yes
Despite subtleties in determining the SnV- Hamiltonian and angular misalignment, it is clear that the readout fidelity reported in this work uses an SnV- center which is only somewhat aligned with the external $\vec{B}$ field, has a spin polarization time of approximately $20\,\mathrm{\mu s}$, and has a cyclicity of approximately $2\times 10^{3}$. At this operating condition, we achieve single-shot readout of the SnV- center’s electronic spin with an average of up to approximately $4$ photons collected during a $50\,\mathrm{\mu s}$ integration window. In Table 1, we compare these results to prior examples of single-shot spin readout of color centers in diamond.
Looking to the future, the misalignment and modest collection efficiency of our experiment emphasize how much SnV- readout can be improved. Even with these nonidealities, the rate of photon collection in our work is much greater than that in the first demonstration of single-shot readout of an SiV- electron spin [15], which reported an average of approximately $6$ photons collected during a 20 ms readout window. This may reflect the SnV-’s higher quantum efficiency $\eta_{q}$: Ref. [44] calculates $\eta_{q}=0.14$ for the SiV- and $\eta_{q}=0.91$ for the SnV-. Subsequent work on SiV- readout has shown greater collection efficiency over faster timescales by optimized use of nanophotonics [19, 20, 21, 22]. Our results here suggest such improvement will also be possible using the SnV-.
We also compare our results to Ref. [43], to our knowledge the only prior published single-shot readout of an SnV- electronic spin. Reference [43] demonstrates readout of $74\%$ fidelity (max of $1.2$ mean counts per shot) collected during a $200\,\mathrm{\mu s}$ readout window. This result is achieved in a bulk sample on an unstrained SnV-. Lack of strain yields high cyclicity with reduced sensitivity to alignment of the magnetic field but also suppresses the magnetic dipole transition necessary for spin control [30]. Finally, we note that Ref. [36] measures single-shot readout of a nuclear spin via a waveguide-integrated SnV-, obtaining a single-shot nuclear spin readout fidelity of $80\%$ in $15\,\mathrm{\mu s}$. In comparison, our work demonstrates a regime in which both single-shot readout of an SnV-’s electronic spin and rapid microwave control of this spin are possible.
@@ C.1 Setup
The setup used in this experiment, Fig. 10, is similar to that in Ref. [30] but optimized for higher transmission on the collection path. The repump (532 nm) and resonant (M-Squared) lasers are both modulated by acousto-optic modulators (4C2C-532-AOM and 4C2C-633-AOM, Gooch and Housego). Typically, we operate with the excitation laser detuned by several GHz from the SnV-’s optical transitions and drive these transitions using the first sideband created by an electro-optic modulator (PM-0S5-PFU-PFU-620, Eospace). By combining two microwave signal generators (SG396, SRS) using a microwave power combiner, we can drive the A1 transition only, the B2 transition only, or both simultaneously as is necessary for implementing charge resonance checks. Pulsing is controlled by an arbitrary waveform generator (Pulse Streamer 8/2, Swabian).
For spin control, microwaves are delivered by a wire bond draped over the diamond chip as in Ref. [30], and using a similar microwave power ($48\,\mathrm{dBm}$ into the cryostat). The wire bond has been replaced from that used in Ref. [30]; we attribute the slower Rabi frequency here to the new wire bond simply being further from the spin.
To focus excitation light and collect signal confocally, a cryogenic objective (LT-APO/VISIR/0.82, Attocube) is placed directly above the chip. This objective has a focal length of 2.87 mm, a numerical aperture of 0.82, and a working distance of 0.65 mm. Signal is then routed out of the cryostat to a 90:10 beam splitter, whose 90% port leads to a free space to fiber coupler. This coupler is connected by a multimode optical fiber (FG025LJA, core diameter $25\pm 3\,\mathrm{\mu m}$, Thorlabs) to a single-photon counting module (SPCM-AQR-15-FC, Perkins Elmer). Detection events are time tagged (Picoharp300, PicoQuant).
Before the fiber coupler, we place a 635 nm long pass filter (635 nm Longwave EdgeBasic, Semrock) to collect the SnV-’s phonon sideband (PSB) only, thus filtering most excitation scatter and also the SnV-’s zero-phonon line. Note that this wavelength still filters a fraction of the PSB; a different filter closer to the zero-phonon line would have lead to greater counts, e.g., a filter at 625 nm. Finally, a 532 nm notch filter (532 nm StopLine, Semrock) is also placed on the collection path to mitigate noise from the repump laser; even though this laser is toggled off during readout, there is some nonzero feedthrough of the AOM used to toggle the green laser which can scatter into the collection path, and so this notch filter is still useful.
@@ C.2 Sources of loss
Source Efficiency Reference or method Path to improve
Quantum efficiency $\eta_{q}$ $80\%-90\%$ [11, 44]
@@
Debye-Waller factor $\eta_{\mathrm{psb}}$ $43\%$ [27] Resonant collection (cross-polarization)
@@
Scattering from chip $\eta_{\mathrm{phot}}$ $5\%$ Simulation, Fig. 11 Nanopillars, waveguides, cavity-QED
Optical path loss $\eta_{\mathrm{path}}$ $35\%$ Transmission, at 670 nm Optimized optics
Detector $\eta_{\mathrm{det}}$ $65\%$ Datasheet spec., at 650 nm Superconducting detectors
Total $\eta$ $0.4\%$ $\eta=\eta_{q}\eta_{\mathrm{psb}}\eta_{\mathrm{phot}}\eta_{\mathrm{path}}\eta_{\mathrm{det}}$ All of the above
@@
Figure 11: (a) Device. The $\langle 100\rangle$ diamond chip contains arrays of nanophotonic structures, both pillars and mesas. Structures have a height of $1\,\mathrm{\mu m}$, and Sn atoms are implanted at a depth of approximately $90\pm 20\,\mathrm{nm}$ [30]. (b) Scan to optimize focal position. The center of the measured lobe (which originates from imaging with a multimode fiber under slightly defocused conditions) corresponds to focal alignment at the edge of a mesa. A scan range of 0.2 V corresponds to a lateral distance of approximately $1\,\mathrm{\mu m}$. (c) Numerical simulation of the far-field scattering pattern from an SnV- center in a mesa. (d) Numerical simulation of 645 nm light collected from both a mesa and (e) a pillar. Simulations include the fraction of light scattered upward, the modal overlap of the objective and the scattering pattern, and their product, which gives the overall fraction of light routed to the collection path. From the simulation in (d), we conclude $\eta_{\mathrm{phot}}\approx 5\%$ of emission enters the collection path. In comparison, similar simulations show a scattering efficiency of $3\%$ from an emitter in bulk diamond at a depth of 100 nm.
@@
In Table 2, we estimate the sources of inefficiency and optical loss expected for this work. Multiplying all sources of inefficiency yields $\eta=\eta_{q}\eta_{\mathrm{psb}}\eta_{\mathrm{phot}}\eta_{\mathrm{col}}\eta_{\mathrm{det}}\approx 0.4\%$. This qualitatively matches the measured values of $\eta\approx 0.2\%$ (Fig. 3) and $\eta\approx 0.1\%$ (Fig. 4).
@@
The dominant source of loss is internal reflection of light into the diamond sample, Fig. 11. By fine-tuning the emitter position using piezoelectric nanopositioners, we find this emitter to actually be in a mesa structure, rather than a nanopillar as stated in Ref. [30]. For an SnV- center implanted at a depth of 100 nm and for structures $1\,\mathrm{\mu m}$ in height, simulations determine scattering from bulk diamond to be approximately $3\%$, scattering from a mesa is approximately $5\%-9\%$, Fig. 11d, and scattering from a pillar is approximately $7\%-28\%$, Fig. 11e. These values are reported at 645 nm, near the maximum of the SnV-’s phonon sideband (PSB). Given the proximity of the color center to a pillar, we use values for an SnV- at the edge of the mesa structure and estimate the (average across the entire PSB) fraction $\eta_{\mathrm{phot}}\approx 5\%$ of emitted light is scattered into our collection path. This number accounts for the modal mismatch between the far field for the SnV- emission and the Gaussian mode accepted by the cryo-objective.
@@
Other sources of loss include the emitter’s quantum efficiency (the probability of nonradiative decay), filtering of the SnV-’s ZPL, loss in the optical path, and detector inefficiency. Finally, Table 2 may not capture every source of loss. Note that we use a 635 nm long pass filter to collect the PSB (see Appendix C.1 for details); this removes approximately $14.5\%$ of the PSB light, and a cutoff of approximately $624\,\mathrm{nm}$ would be best to avoid extra loss [11, 25, 26, 27]. Additionally, while we use multimode fiber to maximize photon collection, some imperfection in the coupling could still be present.
@@
Optimized diamond nanophotonics provides a clear path toward high improvement in collection efficiency (e.g. approximately $25\%$ for inverse-designed grating couplers [56], and approximately $90\%$ for tapered fibers combined with 1D photonic crystal cavities [57]). These devices have also been demonstrated in conjunction with stable SnV-’s [32, 33, 34, 36, 38, 37]. The other sources of loss can be improved upon too. For example, loss associated with filtering of the resonant excitation laser can be avoided by filtering by polarization rather than by wavelength (routinely done in the form of cross-polarization detection). Optical loss in the collection path can be mitigated with higher-quality optics, e.g., by using higher-reflectivity mirrors, and by replacing the 90:10 beam splitter on the collection path with a more transmissive component. Detector inefficiency can be improved by switching to state-of-the-art superconducting detectors.
@@
With improved nanophotonics only, we expect a collection efficiency of order ×$\times 10$ or more, and with all possible improvements we expect an efficiency of order ×$\times 100$ or more can be achieved in future experiments. Even modest improvements in efficiency will allow for much higher single-shot readout fidelity (since fidelity is non-linear with number of photons), operation at lower drive power (see Fig. 4c), faster readout, and single-shot readout at external magnetic fields with greater misalignment.